# Effect of spin in binary neutron star mergers

### v1

@@ Abstract
We investigate the effect of spin on equal and unequal mass binary neutron star mergers using finite-temperature, composition-dependent Hempel-Schaffner-Bielich (SFHo) equation of state, via 3+1 general relativistic hydrodynamics simulations which take into account neutrino emission and absorption. Equal mass cases that have a mass of $M_{1,2}$ =$1.27$, $1.52$ and $2.05M_{\odot}$, result in a supramassive neutron star, a delayed, and a prompt collapse to a black hole, respectively. For all cases, we analyse the effect of initial spin on dynamics, on the structure of the final remnant, its spin evolution, the amount and composition of the ejected matter, gravitational waves, neutrino energy and luminosities. We show that in equal mass binary neutron star mergers, the ejected mass could reach $\sim 0.085M_{\odot}$ for highly aligned-spins ($\chi=0.67$). The black hole which results from such highly spinning, high-mass binary neutron star merger reaches a dimensionless spin of $0.92$; this is the highest spin reached in binary neutron star mergers, to date.
@@ 1 Introduction
The detection of gravitational wave event GW170817 Abbott et al. (2017a) by the Advanced LIGO The LIGO Scientific Collaboration (2015) and Advanced Virgo Acernese et al. (2015) detectors was found to correspond to the inspiral part of a binary neutron star (BNS) merger. The associated electromagnetic counterpart, associated to the short Gamma-Ray Burst (sGRB) event GRB 170817A, which was observed by The Fermi Gamma-Ray Burst Monitor Abbott et al. (2017b), happened $\sim 1.7$s later. This event both questioned a wide range of nuclear and astrophysical theories and showed the long-anticipated link between BNS mergers and sGRBs Eichler et al. (1989); Piran (1992); Ruffert et al. (1995), Abbott et al. (2017b). It constrained individual spins, and constrained (and ruled out) very soft equations of state (EOSs) $(M_{max}<1.97M_{\odot})$ The LIGO Scientific Collaboration and the Virgo Collaboration et al. (2018); Abbott et al. (2019a). It also showed the contribution of BNS mergers to r-process nucleosynthesis and the kilonova Kasen et al. (2017); Smartt et al. (2017); Pian et al. (2017); Troja et al. (2017); Hallinan et al. (2017), the robustness of the general theory of relativity in the strong gravity regime Abbott et al. (2019b) and placed constraints on alternative theories of gravity Sakstein & Jain (2017); Baker et al. (2017). The current detectors are only sensitive to the inspiral period of these events, which allows to measure three important parameters of the system, namely chirp mass Kafka (1988); Finn & Chernoff (1993); Cutler et al. (1993); Cutler & Flanagan (1994), effective spin $\chi$ and tidal deformability $\lambda$, Hinderer (2008); Flanagan & Hinderer (2008); Hinderer et al. (2010); Read et al. (2013); Abbott et al. (2017b). Assuming high spin rates, both aligned with the orbital angular momentum, $(\left|\chi\right|\leq 0.89)$, the chirp mass, effective spin and dimensionless tidal deformability parameter of the binary ($\Lambda$) were measured for the GW170817 event to be $M_{chirp}=1.188^{+0.004}_{-0.002}M_{\odot}$, $\chi\in(0.0,0.09)$, and $\Lambda<800$ (for $90$ per cent credibility level) Abbott et al. (2017a), respectively. Measurement of the tidal deformability parameter is the key to extract the EOS information from gravitational waves.
@@
Creating spinning BNS initial data (ID) was first discussed by Marronetti & Shapiro (2003) who use the conformal thin-sandwich (CTS) method. Tichy (2011, 2012),Tichy (2006, 2009a, 2009b); Tichy et al. (2019), make use of the CTS method and split the fluid velocity into irrotational and rotational parts, the so-called constant rotational velocity (CRV) approach, to create BNS ID. Tsatsin & Marronetti (2013) use the Rapidly Rotating Neutron Star (RNS) code to create BNS ID by mapping two isolated solutions into a single frame. Papenfort et al. (2021) use Fuka which is an extension of the Kadath library Grandclement (2010) and the extended-CTS (xCTS) method to create BNS ID.
@@
We focus on the effect of varying spin on mergers of equal and unequal mass BNS in quasi-circular orbit. The centre-to-centre separation is $40$km. We use the finite temperature, composition dependent Hempel-Schaffner-Bielich (SFHo) EOS by Steiner et al. (2013), taking into account neutrino emission & absorption through 3+1 general relativistic hydrodynamic (GRHD) simulations. The properties of the initial data (ID) are given in Table 1.
@@ 2 Numerical Setup
GRHD evolution is performed by WhiskyTHC Radice & Rezzolla (2012); Radice et al. (2014a, b, 2015). This was built on top of Cactus Allen et al. (1999); Goodale et al. (2003); developers (a, b) and uses the Carpet adaptive mesh refinement (AMR) driver Schnetter et al. (2004) of Cactus. WhiskyTHC is a fourth-order finite difference/finite volume High-Resolution Shock-Capturing (HRSC) code that implements the Valencia formulation Banyuls et al. (1997) of GRHD. It models the neutron stars as perfect fluids with the energy-momentum tensor given as Radice & Rezzolla (2012); Radice et al. (2014a, b, 2015):
@@
We use the finite-temperature, composition-dependent SFHo EOS, which is soft and fully hadronic. It is available on https://stellarcollapse.org, O’Connor & Ott (2010).
Spacetime evolution is performed with CTGamma Pollney et al. (2011), which is based on the publicly available software platform Einstein Toolkit https://einsteintoolkit.org, Löffler et al. (2012); Zilhão & Löffler (2013); Babiuc-Hamilton et al. (2019); Zlochower et al. (2022). We use the constraint damping z4c formulation by Bernuzzi & Hilditch (2010), which is used within CTGamma, with moving puncture gauge conditions.
@@ 3 Results
All post-processing is performed using PostCactus Kastaun (2021). Since the simulations with $M_{tot}=2.55M_{\odot}$ are performed with two resolutions, the results given in Tables 1, 2, LABEL:ejecta_properties_table, LABEL:bh_properties, LABEL:neutrino_table, LABEL:peak_neutrino_lum, LABEL:gw_energy and LABEL:gw_angmom that have a "$\pm$" show the difference between the two resolutions. All figures are created using data from models with resolution $h=307$m, for $M_{tot}=2.55M_{\odot}$.
@@ 3.1 Using two versions of the Einstein Toolkit
To show the consistency between the "Mayer" and "Sophie Kowalevski" release of the Einstein Toolkit, see section 1 and section 2 for details, we compared irrotational cases and found the difference in $M_{ej}$, the mass weighted average electron fraction, $\langle Y_{e}\rangle$, and velocity, $\langle v\rangle$. to be less than $<10^{-8}$. The difference in release of both the total energy and angular momentum (in $+z$ direction) due to gravitational radiation is less than $<4\times 10^{-5}$. This comparison shows that the version difference does not have any impact for our study.
@@ 3.5.1 Energy and angular momentum release
The aligned spinning cases lead to higher angular momentum loss than the irrotational case for $M_{tot}=3.05$, except for the unequal mass cases and $4.10M_{\odot}$. The maximum angular momentum is radiated by the irrotational case for $M_{tot}=2.55M_{\odot}$. For this total mass, aligned spinning cases have higher angular momentum loss than the anti-aligned spinning cases.
@@ 3.5.2 Frequency and spectra
We show the effect of spin on the instantanous frequency at the instant merger $f_{merger}$ and the post-merger peak frequency $f^{peak}_{l,m}$, for different modes. We present the merger and the peak frequencies in Table 2. In this table we only provide the instantaneous merger frequencies for the binaries that promptly collapse to the BH and $f^{peak}_{2,0}$ for $\chi=0.67$ at the highest resolution because the difference between the two resolution is greater than $100$ per cent. The gravitational wave spectra are highly sensitive to numerical resolution. As the peak frequency for the $(2,0)$ mode does not converge between the two resolutions, we assume the highest resolution to be closer to the actual value.
The merger and peak frequency for $(2,1)$ mode increases up to $\sim 1.5$kHz, and $\sim 1$kHz, for $\chi=-0.65$, compared to the irrotational case, respectively. The change in peak frequencies of the $(2,2)$ and $(3,3)$ modes are at most $\sim 0.4$kHz for the $\chi=-0.65$ cases. The shift in the gravitational wave frequency of the $(2,0)$ mode increases by $\sim 4$kHz for the $\chi=0.2$ case as comapred to the irrotational case.
@@ 3.6 Ejecta properties
We calculate the ejected matter properties, unbound matter that does not fall back, from a surface placed at $\sim 443$km. We use the Outflow thorn with the geodesic criteria Kastaun & Galeazzi (2015) $u_{t}<-1$, where $u_{t}$ is the time component of the four-velocity. The ejected matter properties are presented in Table LABEL:ejecta_properties_table. Time calculations are performed in the frame of reference of the binary. Electron fraction and velocity of the matter are given as mass-weighted averages. The presence of fast-moving ejecta, which we identify to be ejecta moving with mass weighted velocity equal to or greater than $0.6$c (Radice et al. (2018) and references therein) is presented for number of cases. The total mass of the ejected matter increases for the all considered spin magnitudes and orientations, except for many of the $M_{tot}=4.10M_{\odot}$ models. It reaches $\sim 0.085M_{\odot}$ for the extreme aligned spin case $\chi=0.67$.
@@
Table 3: The properties of ejected matter for all models at two dynamical times $t_{dyn,1}$ and $t_{dyn,2}$ in the frame of reference of the binary (see section 3 for the definition of these times). The ejecta properties are extracted at a surface located at a distance $r=443$km. $M_{ej}$ and $M_{ej}\geq 0.6$c show the total mass of the ejecta and the fast-moving ejecta (Radice et al. (2018) and references therein), $\langle v\rangle$ and $\langle Y_{e}\rangle$ represent the mass-weighted average velocity, and electron fraction of the ejecta, respectively. The values given with $\pm$ show the difference between two resolutions.
Model $M_{\text{ej}}$ ($10^{-3}M_{\odot}$) $M_{\text{ej}}\geq 0.6c$ ($10^{-5}M_{\odot}$) $\langle Y_{e}\rangle$ $\langle v\rangle$ ($c$)
$t_{dyn,1}$ $t_{dyn,2}$ $t_{dyn,1}$ $t_{dyn,2}$ $t_{dyn,1}$ $t_{dyn,2}$ $t_{dyn,1}$ $t_{dyn,2}$
$\chi=0[1]$ $0.46\pm 0.09$ 0.00 - $0.27\pm 0.02$ $0.31\pm 0.03$
$\chi=-0.2$ $2.06\pm 0.24$ $3.40\pm 1.00$ $1.30\pm 1.20$ $1.30\pm 1.20$ $0.18\pm 0.00$ $0.18\pm 0.01$ $0.32\pm 0.01$ $0.26\pm 0.02$
$\chi=0.2$ $1.20\pm 0.62$ 0.00 - $0.10\pm 0.05$ $0.23\pm 0.00$
$\chi=-0.4$ $3.32\pm 1.19$ $6.96\pm 2.62$ $1.42\pm 0.48$ $1.42\pm 0.48$ $0.26\pm 0.00$ $0.26\pm 0.01$ $0.35\pm 0.04$ $0.27\pm 0.02$
$\chi=0[2]$ $0.75\pm 0.01$ $2.35\pm 0.97$ 0.00 0.00 $0.17\pm 0.02$ $0.20\pm 0.04$ $0.25\pm 0.00$ $0.20\pm 0.01$
$\chi=0.4$ $9.43\pm 0.33$ $24.29\pm 2.02$ 0.00 0.00 $0.06\pm 0.00$ $0.06\pm 0.00$ $0.25\pm 0.00$ $0.20\pm 0.01$
$\chi=-0.65$ $12.03\pm 0.86$ $15.57\pm 0.48$ $26.00\pm 2.86$ $26.00\pm 2.86$ $0.21\pm 0.02$ $0.20\pm 0.01$ $0.31\pm 0.02$ $0.28\pm 0.01$
$\chi=0[3]$ $48.82\pm 4.29$ $75.58\pm 3.73$ $0.01\pm 0.01$ $0.01\pm 0.01$ $0.05\pm 0.00$ $0.06\pm 0.00$ $0.26\pm 0.01$ $0.23\pm 0.00$
$\chi=0.67$ $55.77\pm 8.83$ $85.37\pm 1.58$ 0.00 0.00 $0.04\pm 0.01$ $0.05\pm 0.01$ $0.25\pm 0.01$ $0.22\pm 0.00$
$305q1\chi 0$ 0.064 - 0.00 - 0.31 - 0.31 -
$305q1\chi 0.4$ 7.644 10.541 0.00 0.00 0.09 0.08 0.21 0.20
$305q2.05\chi 0$ 0.009 - 0.00 - 0.29 - 0.26 -
$305q2.05\chi 0.4$ 3.409 33.826 0.37 0.42 0.08 0.07 0.29 0.20
$41q1\chi 0[1]$ 1.500 - 18.07 - 0.36 - 0.46 -
$0.6541q1\chi-$ 0.65 0.003 - 0.02 - 0.27 - 0.35 -
$41q1\chi 0[2]$ 0.369 - 3.61 - 0.15 - 0.46 -
$41q1\chi 0.67$ 3.948 - 0.00 - 0.10 - 0.40 -
@@
At time $t_{dyn,1}$, a increase of the mass ratio decreases the total mass of the ejected matter for both the irrotational and the spinning cases.

### v2

@@ Abstract
We investigate the effect of spin on equal and unequal mass binary neutron star mergers using finite-temperature, composition-dependent Hempel-Schaffner-Bielich (SFHo) equation of state, via 3+1 general relativistic hydrodynamics simulations which take into account neutrino emission and absorption. Equal mass cases that have a mass of $M_{1,2}$ =$1.27$, $1.52$ and $2.05M_{\odot}$, result in a supramassive neutron star, a delayed, and a prompt collapse to a black hole, respectively. For all cases, we analyse the effect of initial spin on dynamics, on the structure of the final remnant, its spin evolution, the amount and composition of the ejected matter, gravitational waves, neutrino energy and luminosities, and disc masses. We show that in equal mass binary neutron star mergers, the ejected mass could reach $\sim 0.085M_{\odot}$ for highly aligned-spins ($\chi=0.67$). The black hole which results from such highly spinning, high-mass binary neutron star merger reaches a dimensionless spin of $0.92$; this is the highest spin reached in binary neutron star mergers, to date.
@@ 1 Introduction
The detection of gravitational wave event GW170817 Abbott et al. (2017a) by the Advanced LIGO The LIGO Scientific Collaboration (2015) and Advanced Virgo Acernese et al. (2015) detectors was found to correspond to the inspiral part of a binary neutron star (BNS) merger. The associated electromagnetic counterpart, associated to the short Gamma-Ray Burst (sGRB) event GRB 170817A, which was observed by The Fermi Gamma-Ray Burst Monitor Abbott et al. (2017b), happened $\sim 1.7$s later. This event both questioned a wide range of nuclear and astrophysical theories and showed the long-anticipated link between BNS mergers and sGRBs Eichler et al. (1989); Piran (1992); Ruffert et al. (1995, 1997); Ruffert & Janka (2001), Abbott et al. (2017b). It constrained individual spins, and constrained (and ruled out) very soft equations of state (EOSs) $(M_{max}<1.97M_{\odot})$ The LIGO Scientific Collaboration and the Virgo Collaboration et al. (2018); Abbott et al. (2019a). It also showed the contribution of BNS mergers to r-process nucleosynthesis and the kilonova Kasen et al. (2017); Smartt et al. (2017); Pian et al. (2017); Troja et al. (2017); Hallinan et al. (2017), the robustness of the general theory of relativity in the strong gravity regime Abbott et al. (2019b) and placed constraints on alternative theories of gravity Sakstein & Jain (2017); Baker et al. (2017). The current detectors are only sensitive to the inspiral period of these events, which allows to measure three important parameters of the system, namely chirp mass Kafka (1988); Finn & Chernoff (1993); Cutler et al. (1993); Cutler & Flanagan (1994), effective spin $\chi$ and tidal deformability $\lambda$, Hinderer (2008); Flanagan & Hinderer (2008); Hinderer et al. (2010); Read et al. (2013); Abbott et al. (2017b). Assuming high spin rates, both aligned with the orbital angular momentum, $(\left|\chi\right|\leq 0.89)$, the chirp mass, effective spin and dimensionless tidal deformability parameter of the binary ($\Lambda$) were measured for the GW170817 event to be $M_{chirp}=1.188^{+0.004}_{-0.002}M_{\odot}$, $\chi\in(0.0,0.09)$, and $\Lambda<800$ (for $90$ per cent credibility level) Abbott et al. (2017a), respectively. Measurement of the tidal deformability parameter is the key to extract the EOS information from gravitational waves.
@@
Creating spinning BNS initial data (ID) was first discussed by Marronetti & Shapiro (2003) who use the conformal thin-sandwich (CTS) method. Tichy (2006, 2009a, 2009b, 2011, 2012); Tichy et al. (2019), make use of the CTS method and split the fluid velocity into irrotational and rotational parts, the so-called constant rotational velocity (CRV) approach, to create BNS ID. Tsatsin & Marronetti (2013) use the Rapidly Rotating Neutron Star (RNS) code to create BNS ID by mapping two isolated solutions into a single frame. Tacik et al. (2015) use the extended-CTS (xCTS) method and the CRV approach of Tichy to create BNS ID. Papenfort et al. (2021) use Fuka which is an extension of the Kadath library Grandclement (2010) and the xCTS method to create BNS ID.
@@
We focus on the effect of varying spin on mergers of equal and unequal mass BNS in quasi-circular orbit. The centre-to-centre separation is $40$km. We use the finite-temperature, composition-dependent Hempel-Schaffner-Bielich (SFHo) EOS by Steiner et al. (2013), taking into account neutrino emission & absorption through 3+1 general relativistic hydrodynamic (GRHD) simulations. The properties of the initial data (ID) are given in Table 1.
@@ 2 Numerical Setup
GRHD evolution is performed by WhiskyTHC Radice & Rezzolla (2012); Radice et al. (2014a, b, 2015). This was built on top of Cactus Allen et al. (1999); Goodale et al. (2003); Cactus developers (2023a, b) and uses the Carpet adaptive mesh refinement (AMR) driver Schnetter et al. (2004) of Cactus. WhiskyTHC is a fourth-order finite difference/finite volume High-Resolution Shock-Capturing (HRSC) code that implements the Valencia formulation Banyuls et al. (1997) of GRHD. It models the neutron stars as perfect fluids with the energy-momentum tensor given as Radice & Rezzolla (2012); Radice et al. (2014a, b, 2015):
@@
We use the finite-temperature, composition-dependent SFHo EOS, which is soft and fully hadronic. It is available on Stellarcollapse (2017), O’Connor & Ott (2010).
Spacetime evolution is performed with CTGamma Pollney et al. (2011), which is based on the publicly available software platform Einstein Toolkit Löffler et al. (2012); Zilhão & Löffler (2013); Babiuc-Hamilton et al. (2019); Zlochower et al. (2022); The Einstein Toolkit (2024). We use the constraint damping z4c formulation by Bernuzzi & Hilditch (2010), which is used within CTGamma, with moving puncture gauge conditions.
@@ 3 Results
All post-processing is performed using PostCactus  Kastaun (2021) and Scidata Radice (2023). Since the simulations with $M_{tot}=2.55M_{\odot}$ are performed with two resolutions, the results given in Tables 1, 2, LABEL:ejecta_properties_table, LABEL:bh_properties, LABEL:neutrino_table, LABEL:peak_neutrino_lum, LABEL:gw_energy and LABEL:gw_angmom that have a "$\pm$" show the difference between the two resolutions. All figures are created using data from models with resolution $h=307$m, for $M_{tot}=2.55M_{\odot}$.
@@ 3.1 Using two versions of the Einstein Toolkit
To show the consistency between the "Mayer" and "Sophie Kowalevski" release of the Einstein Toolkit, see section 1 and section 2 for details, we compared irrotational cases and found the difference in $M_{ej}$, the mass weighted average electron fraction, $\langle Y_{e}\rangle$, and velocity, $\langle v\rangle$, to be less than $<10^{-8}$. The difference in release of both the total energy and angular momentum (in $+z$ direction) due to gravitational radiation is less than $<4\times 10^{-5}$. This comparison shows that the version difference does not have any impact for our study.
@@ 3.5.1 Energy and angular momentum release
The aligned spinning cases lead to higher angular momentum loss than the irrotational case for $M_{tot}=3.05M_{\odot}$, except for the unequal mass cases, and $4.10M_{\odot}$. The maximum angular momentum is radiated by the irrotational case for $M_{tot}=2.55M_{\odot}$. For this total mass, aligned spinning cases have higher angular momentum loss than the anti-aligned spinning cases.
@@ 3.5.2 Frequency and spectra
We show the effect of spin on the instantaneous frequency at the instant merger $f_{merger}$ and the post-merger peak frequency $f^{peak}_{l,m}$, for different modes. We present the merger and the peak frequencies in Table 2. In this table we only provide the instantaneous merger frequencies for the binaries that promptly collapse to the BH and $f^{peak}_{2,0}$ for $\chi=0.67$ at the highest resolution because the difference between the two resolution is greater than $100$ per cent. The gravitational wave spectra are highly sensitive to numerical resolution. As the peak frequency for the $(2,0)$ mode does not converge between the two resolutions, we assume the highest resolution to be closer to the actual value.
The merger and peak frequency for $(2,1)$ mode increases up to $\sim 1.5$kHz, and $\sim 1$kHz, for $\chi=-0.65$, compared to the irrotational case, respectively. The change in peak frequencies of the $(2,2)$ and $(3,3)$ modes are at most $\sim 0.4$kHz for the $\chi=-0.65$ cases. The shift in the gravitational wave frequency of the $(2,0)$ mode increases by $\sim 4$kHz for the $\chi=0.2$ case as compared to the irrotational case.
@@ 3.6 Ejecta properties
We calculate the ejected matter properties, unbound matter that does not fall back, from a surface placed at $\sim 443$km. We use the Outflow thorn with the geodesic criteria Kastaun & Galeazzi (2015) $u_{t}<-1$, where $u_{t}$ is the time component of the four-velocity. The ejected matter properties are presented in Table LABEL:ejecta_properties_table. Time calculations are performed in the frame of reference of the binary. Electron fraction and velocity of the matter are given as mass-weighted averages. The presence of fast-moving ejecta, which we identify to be ejecta moving with mass weighted velocity equal to or greater than $0.6$c (Radice et al. (2018) and references therein) is presented for a number of cases. The total mass of the ejected matter increases for the all considered spin magnitudes and orientations, except for many of the $M_{tot}=4.10M_{\odot}$ models. It reaches $\sim 0.085M_{\odot}$ for the extreme aligned spin case $\chi=0.67$.
@@
Table 3: The properties of ejected matter for all models at two dynamical times $t_{dyn,1}$ and $t_{dyn,2}$ in the frame of reference of the binary, and disc masses at $t_{dyn,2}$ if extant (see section 3 for the definition of these times). The ejecta properties are extracted at a surface located at a distance $r=443$km. $M_{ej}$ and $M_{ej}\geq 0.6$c show the total mass of the ejecta and the fast-moving ejecta (Radice et al. (2018) and references therein), $M_{disc}$, $\langle v\rangle$ and $\langle Y_{e}\rangle$ represent the disc mass, mass-weighted average velocity, and electron fraction of the ejecta, respectively. We provide disc masses only for a resolution of $h=307$m for binaries with $M_{tot}=2.55M_{\odot}$.The values given with $\pm$ show the difference between two resolutions.
Model $M_{\text{ej}}$ ($10^{-3}M_{\odot}$) $M_{\text{ej}}\geq 0.6c$ ($10^{-5}M_{\odot}$) $M_{\text{disc}}$ $\langle Y_{e}\rangle$ $\langle v\rangle$ ($c$)
$t_{dyn,1}$ $t_{dyn,2}$ $t_{dyn,1}$ $t_{dyn,2}$ ($10^{-3}M_{\odot}$) $t_{dyn,1}$ $t_{dyn,2}$ $t_{dyn,1}$ $t_{dyn,2}$
$\chi=0[1]$ $0.46\pm 0.09$ 0.00 - $127$ $0.27\pm 0.02$ $0.31\pm 0.03$
$\chi=-0.2$ $2.06\pm 0.24$ $3.40\pm 1.00$ $1.30\pm 1.20$ $1.30\pm 1.20$ $198$ $0.18\pm 0.00$ $0.18\pm 0.01$ $0.32\pm 0.01$ $0.26\pm 0.02$
$\chi=0.2$ $1.20\pm 0.62$ 0.00 - $164$ $0.10\pm 0.05$ $0.23\pm 0.00$
$\chi=-0.4$ $3.32\pm 1.19$ $6.96\pm 2.62$ $1.42\pm 0.48$ $1.42\pm 0.48$ $130$ $0.26\pm 0.00$ $0.26\pm 0.01$ $0.35\pm 0.04$ $0.27\pm 0.02$
$\chi=0[2]$ $0.75\pm 0.01$ $2.35\pm 0.97$ 0.00 0.00 $134$ $0.17\pm 0.02$ $0.20\pm 0.04$ $0.25\pm 0.00$ $0.20\pm 0.01$
$\chi=0.4$ $9.43\pm 0.33$ $24.29\pm 2.02$ 0.00 0.00 $365$ $0.06\pm 0.00$ $0.06\pm 0.00$ $0.25\pm 0.00$ $0.20\pm 0.01$
$\chi=-0.65$ $12.03\pm 0.86$ $15.57\pm 0.48$ $26.00\pm 2.86$ $26.00\pm 2.86$ $251$ $0.21\pm 0.02$ $0.20\pm 0.01$ $0.31\pm 0.02$ $0.28\pm 0.01$
$\chi=0[3]$ $48.82\pm 4.29$ $75.58\pm 3.73$ $0.01\pm 0.01$ $0.01\pm 0.01$ $221$ $0.05\pm 0.00$ $0.06\pm 0.00$ $0.26\pm 0.01$ $0.23\pm 0.00$
$\chi=0.67$ $55.77\pm 8.83$ $85.37\pm 1.58$ 0.00 0.00 $304$ $0.04\pm 0.01$ $0.05\pm 0.01$ $0.25\pm 0.01$ $0.22\pm 0.00$
$305q1\chi 0$ 0.064 - 0.00 - $0.839$ 0.31 - 0.31 -
$305q1\chi 0.4$ 7.644 10.541 0.00 0.00 $274$ 0.09 0.08 0.21 0.20
$305q2.05\chi 0$ 0.009 - 0.00 - $179$ 0.29 - 0.26 -
$305q2.05\chi 0.4$ 3.409 33.826 0.37 0.42 $247$ 0.08 0.07 0.29 0.20
$41q1\chi 0[1]$ 1.500 - 18.07 - $1.00$ 0.36 - 0.46 -
$0.6541q1\chi-$ 0.65 0.003 - 0.02 - $0.006$ 0.27 - 0.35 -
$41q1\chi 0[2]$ 0.369 - 3.61 - $145$ 0.15 - 0.46 -
$41q1\chi 0.67$ 3.948 - 0.00 - $363$ 0.10 - 0.40 -
@@
At time $t_{dyn,1}$, an increase of the mass ratio decreases the total mass of the ejected matter for both the irrotational and the spinning cases.